# Kudla Millson lift of toric cycles and restriction of Hilbert modular forms

### v3

@@ 3.5.  Computations of the Hilbert modular form
$\varphi^{E}_{\operatorname{KM}}={\left.\kern-1.2pt\varphi^{V}_{\operatorname{KM}}\vphantom{\big{|}}\right|_{\mathcal{D}^{+}_{T}}}\in\Omega^{q}(\mathcal{D}^{+}_{T}).$ (3.112)
Suppose that $E$ is a field Since $\operatorname{T}$ is of maximal real rank, $S_{3}$ is empty by Proposition 2.6. Recall that $q=|S_{1}|$ by Proposition 2.6. At a place $\sigma\in S_{2}$, the algebra is $E_{\sigma}\simeq\mathbb{C}$ and the space $\mathcal{D}_{\sigma}^{+}$ is a point. So the Kudla-Millson form is of degree 00 and is simply the Gaussian
@@
At a place $\sigma\in S_{1}$, the algebra is $E_{\sigma}\simeq\mathbb{R}\times\mathbb{R}$ and $\mathcal{D}_{\sigma}^{+}\simeq\mathbb{R}_{>0}$. An element $v$ in $E_{\sigma}$ is sent to $(v_{\sigma},v^{\prime}_{\sigma})$ in $E_{\sigma}$. The Kudla-Millson form (see [Bra23a]) is given by
@@
for $g_{\infty}\in\operatorname{SL}_{2}(F_{\mathbb{R}})$. Note that since $E$ is a field, any nonzero $v\in E$ must satisfy $v_{\sigma}v_{\sigma}^{\prime}\neq 0$ at any place $\sigma\in S_{1}$. Let $\tau\coloneqq(\tau_{\sigma})_{\sigma}\in\mathscr{H}^{q}$ be the image of $(i,\dots,i)$ by $g_{\infty}$ and write
@@
for $m\in F$.
@@
[…]
if $m=\operatorname{N}_{E/F}(v)$ is totally positive, and the integral is 00 otherwise.
@@
[…]
@@
after the substitution $t_{\sigma}=\sqrt{\left|\frac{v_{\sigma}}{v_{\sigma}^{\prime}}\right|}u$. Note that the assumption that $\chi_{\sigma}(t_{\sigma})=\operatorname{sgn}(t_{\sigma})$ is crucial here, since the integral would be 00 otherwise. After substituting $\beta=u^{2}$ we find that this integral is
@@
[…]
@@
$\Theta_{\varphi}(\tau_{1},\cdots,\tau_{d})=\widetilde{\Theta}_{\varphi}(\tau_{1},\cdots,\tau_{d},t)dt^{\times}\in\Omega^{q}(Y_{T}).$ (3.121)
By the (3.116) we have
$\widetilde{\Theta}_{\varphi}(\tau_{1},\cdots,\tau_{d},t)=(y_{1}\cdots y_{d})^{-\frac{1}{2}}\sum_{v\in E}(\omega(g_{\infty},t)\varphi)(v)$ (3.122)
@@
$\Theta_{\varphi}(\tau_{1},\cdots,\tau_{d},\chi)$ $=\frac{2^{q}}{\operatorname{vol}(K_{T})}\int_{E^{1}\backslash E^{1}_{\mathbb{A}}}(y_{1}\cdots y_{d})^{-\frac{1}{2}}\sum_{v\in E}(\omega(g_{\infty},t)\varphi)(v)\chi(t)dt^{\times}$
[…]
@@
Let $\chi$ be a character on a anisotropic maximal $\mathbb{Q}$-torus $\operatorname{T}$ of maximal real rank. Then $\Theta_{\varphi}(\tau,C_{\chi})$ is a cusp form.
@@
Suppose that $F$ is a field, and $E$ a quadratic field extension. Then the only $v\in E$ for which $Q_{\alpha}(v,v)=\operatorname{N}_{E/F}(v,v)=0$ is $v=0$. Since $v_{\sigma}=v_{\sigma}^{\prime}=0$, we see from (3.126) that the constant term $c^{(0)}_{\varphi}(C_{\chi})$ is zero.
@@ 3.6.  Generating series of intersection numbers
Let $(V,Q)$ be a quadratic space of even dimension $2d$ over $\mathbb{Q}$. Let $L$ be an even integral lattice and $\varphi_{f}=\mathbf{1}_{\widehat{L}}\in\mathcal{S}(V_{\widehat{\mathbb{Q}}})$ its characteristic function. We describe the geometric features of the Kudla-Millson form, that allows us to have a nice geometric interpretation of Kudla-Millson lift $\Theta_{\varphi}(\tau,C)$ defined in (3.3.1).
@@
$\mathcal{D}^{+}_{v}\coloneqq\left\{z\in\mathcal{D}^{+}\left|z\subset v^{\perp}\right.\right\}.$ (3.131)
@@
$\mathcal{C}=\operatorname{SO}_{V}(\mathbb{Q})^{+}\backslash\operatorname{SO}_{V}(\widehat{\mathbb{Q}})/K_{f}$ (3.132)
@@
$\Theta_{\varphi}(\tau)=y^{-\frac{p+q}{2}}\sum_{v\in V}(\omega(g_{\tau})\varphi)(v)=\sum_{n=0}^{\infty}\Theta_{\varphi}^{(n)}(\tau)q^{n}$ (3.142)
@@
$\Theta_{\varphi}^{(n)}(\tau)\coloneqq\sum_{\begin{subarray}{c}v\in V\\ Q(v,v)=2n\end{subarray}}\varphi_{0}(\sqrt{y}v)\varphi_{f}(v)=\sum_{\begin{subarray}{c}v\in L\\ Q(v,v)=2n\end{subarray}}\varphi_{0}(\sqrt{y}v).$ (3.147)
The form $\Theta_{\varphi}^{(n)}(\tau)$ is closed, and for positive $n$ it represents a Poincaré dual of the special cycle $C_{n}(\varphi)$ in $H^{q}(Y,\mathbb{C})$. When $V$ is anisotropic, the constant term is $\varphi_{0}(0)\varphi_{f}(0)$ where $\varphi_{0}(0)$ represents the Euler class of the tautological bundle over $\mathcal{D}^{+}$.
@@
$\Theta_{\varphi}(\tau,C)=\int_{C}\Theta_{\varphi}(\tau)$ (3.148)
@@
$\int_{C}\Theta_{\varphi}(\tau)=c^{0}_{\varphi}(C)+\sum_{n=1}^{\infty}\int_{C}\Theta_{\varphi}^{(n)}(\tau)q^{n}=c^{0}_{\varphi}(C)+\sum_{n=1}^{\infty}\langle C,C_{n}(\varphi)\rangle q^{n}.$ (3.149)
When the cycles $C$ and $C_{n}(\varphi)$ intersect transversely, then $\langle C,C_{n}(\varphi)\rangle$ is the signed intersection number
@@
Let $E=E_{1}\times\cdots\times E_{r}$ with fixed subalgebra $F=F_{1}\times\cdots\times F_{r}$. So the torus is a product $\operatorname{T}(\mathbb{Q})=\operatorname{T}_{1}(\mathbb{Q})\times\cdots\times\operatorname{T}_{r}(\mathbb{Q})$ where $\operatorname{T}_{i}(\mathbb{Q})\simeq E_{i}^{1}$. Let $\chi=\chi_{1}\times\cdots\times\chi_{r}$ where $\chi_{i}\colon\operatorname{T}(\mathbb{Q})\backslash\operatorname{T}(\mathbb{A})\longrightarrow\mathbb{C}^{\times}$. The Schwartz function splits as $\varphi_{f}=\sum_{\beta}\varphi_{f}^{\beta,1}\otimes\cdots\otimes\varphi_{f}^{\beta,r}\in\mathcal{S}(\widehat{E}_{1})\otimes\cdots\otimes\mathcal{S}(\widehat{E}_{r})$, as in Proposition 3.2.
@@ 4.  Example of biquadratic fields
We consider the setting of a split quadratic space of signature $(2,2)$. The setting is similar to [Bra23], but we consider the integral over a product of compact geodesics instead of the the product of a compact geodesic with $\gamma_{\infty}$.
@@
The connected component $\operatorname{GSpin}_{V}(\mathbb{R})^{+}=\operatorname{GL}_{2}(\mathbb{R})^{+}\times_{\det}\operatorname{GL}_{2}(\mathbb{R})^{+}$ consists of pairs of matrices with (same) positive determinant. It acts transitively on $\mathscr{H}\times\mathscr{H}$ by Möbius transformations in both factors. Let $\widetilde{K}_{\infty}\subset\operatorname{GSpin}_{V}(\mathbb{R})^{+}$ be the stabilizer of a point in $\mathscr{H}\times\mathscr{H}$, so that ×$\mathscr{H}\times\mathscr{H}\simeq\operatorname{GSpin}_{V}(\mathbb{R})^{+}/\widetilde{K}_{\infty}$. The stabilizer of $(i,i)$ is $\mathbb{R}_{>0}(\operatorname{SO}(2)\times\operatorname{SO}(2))$. We can extend the Weil representation from the pair $\operatorname{SO}_{V}(\mathbb{Q}_{v})\times\operatorname{SL}_{2}(\mathbb{Q}_{v})$ to the pair $\operatorname{GSpin}_{V}(\mathbb{Q}_{v})\times\operatorname{SL}_{2}(\mathbb{Q}_{v})$ by
@@ 4.1.  Locally symmetric space
For the locally symmetric spaces and the special cycles in this setting, we refer to [Bra23] for more details. Let $M_{0}(p)\subset\operatorname{Mat}_{2}(\mathbb{Q})$ be the lattice
@@
that satisfies $\operatorname{Mat}_{2}(\mathbb{Q})^{+}\cap\widehat{M}_{0}(p)=M_{0}(p)$. Let $\varphi_{f}\in\mathcal{S}(\operatorname{Mat}_{2}(\widehat{\mathbb{Z}}))$ be the characteristic function of $\widehat{M}_{0}(p)$. It is preserved by the open compact $\widetilde{K}_{0}(p)$, where $\widetilde{K}_{0}(p)\coloneqq K_{0}(p)\times_{\det}K_{0}(p)$ is an open compact in $\operatorname{GSpin}_{V}(\widehat{\mathbb{Z}})$ and
$K_{0}(p)=\left\{\left.\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{GL}_{2}(\widehat{\mathbb{Z}})\right|\ \in p\mathbb{Z}\right\}.$ (4.22)
@@
by the strong approximation $\operatorname{GL}_{2}(\widehat{\mathbb{Q}})=\operatorname{GL}_{2}(\mathbb{Q})^{+}K_{0}(p)$.
@@
$\mathscr{H}$$\lhook\joinrel\longrightarrow\mathscr{H}\times\mathscr{H}⸦$
@@ 4.2.  Maximal torus in $\operatorname{GSpin}_{V}$
Let $L=\mathbb{Q}(\sqrt{D})$ be a real quadratic field of fundamental discriminant $D$ and suppose that $p$ is split in $L$. Let $\mathcal{O}_{L}$ be the ring of integers. Let us fix an integer $r\in\mathbb{Z}$ such that $r^{2}\equiv D\pmod{4p}$. A form $[a,b,c]=ax^{2}+bxy+cy^{2}$ of squarefree discriminant $D$ is called a Heegner form at $p$ if it satisfies $a\equiv 0\pmod{p}$ and $b\equiv r\pmod{p}$. Let $(u,v)$ be a positive fundamental solution to the Pell equation $u^{2}-Dv^{2}=1$. To every primitive Heegner form $[a,b,c]$ at $p$ of discriminant $D=b^{2}-4ac$ we can associate the matrix
@@
Instead of taking a maximal algebraic $\mathbb{Q}$-torus $\operatorname{T}$ in $\operatorname{SO}_{V}$, let us start with a maximal $\mathbb{Q}$ torus $\widetilde{\operatorname{T}}$ in $\operatorname{GSpin}_{V}$. The image $\operatorname{T}(\mathbb{Q})=\rho(\widetilde{\operatorname{T}}(\mathbb{Q}))$ in $\operatorname{SO}_{V}(\mathbb{Q})$ is a maximal $\mathbb{Q}$-torus and we have an exact sequence
$1\xrightarrow{\ \ \ \ \ }\mathbb{Q}^{\times}\xrightarrow{\ \ \ \ \ }\widetilde{\operatorname{T}}(\mathbb{Q})\xrightarrow{\ \ \rho\ \ }\operatorname{T}(\mathbb{Q})\xrightarrow{\ \ \ \ \ }1$ (4.39)
Let $L_{1}$ and $L_{2}$ be two real quadratic fields with distinct discriminants $D_{1}$ and $D_{2}$. Suppose that $p$ is split in both fields. Let $[N_{1},r_{1},1]$ and $[N_{2},r_{2},1]$ be the two principal [^3] It represents the unit in the narrow class group. Heegner forms where $N_{i}\coloneqq\frac{r_{i}^{2}-D_{i}}{4}$. Let $\phi_{i}\colon L_{i}\lhook\joinrel\longrightarrow\operatorname{Mat}_{2}(\mathbb{Q})$ be the two associated optimal embeddings, given by
@@
We can then map $\operatorname{GSpin}_{V}(\mathbb{Q})$ in $\operatorname{End}_{\mathbb{Q}}(\operatorname{Mat}_{2}(\mathbb{Q}))$ by sending $(g_{1},g_{2})$ to $\det(g_{2})g_{1}\otimes g_{2}$. The map is compatible with the actions of $\operatorname{GSpin}_{V}(\mathbb{Q})$ and $\operatorname{End}_{\mathbb{Q}}(V)$ on $V$
@@
Since $Q(x,y)=\operatorname{Tr}(xy^{*})=\operatorname{Tr}(x^{\ast}y)$, we have $Q(xb,y)=Q(x,yb^{*})$and $Q(ax,y)=Q(x,a^{*}y)$. Hence, the involution is given by $\epsilon_{Q}(a\otimes b)=a^{*}\otimes b^{*}$. The image of $\widetilde{\operatorname{T}}(\mathbb{Q})$ in $\operatorname{End}_{\mathbb{Q}}(\operatorname{Mat}_{2}(\mathbb{Q}))\simeq\operatorname{Mat}_{2}(\mathbb{Q})\otimes_{\mathbb{Q}}\operatorname{Mat}_{2}(\mathbb{Q})$ is
@@
The étale algebra of endomorphisms commuting with $J$ is $E=L_{1}\otimes L_{2}$. Note that when restricted to $L_{i}$ the involution $x^{*}$ acts like the Galois involution $x\mapsto x^{\prime}$, i.e. we have $\phi(x^{\prime})=\phi(x)^{*}$. Hence, the involution $\epsilon_{Q}$ restricted to $E$ is $\epsilon_{Q}(x_{1}\otimes x_{2})=x_{1}^{\prime}\otimes x_{2}^{\prime}$ where $x_{i}^{\prime}$ is the Galois conjugate of $x_{i}$. It follows that the étale algebra is $E\simeq\mathbb{Q}(\sqrt{D_{1}},\sqrt{D_{2}})$ with the involution sending $\sqrt{D_{i}}$ to $-\sqrt{D_{i}}$. ∎
@@
$\mathbb{Q}(\sqrt{D_{1}},\sqrt{D_{2}})^{1}\simeq\operatorname{T}(\mathbb{Q})\simeq\rho\left(L_{1}^{\times}\times_{\operatorname{N}}L_{2}^{\times}\right)$ (4.49)
@@
where $\widetilde{\operatorname{T}}_{0}(\mathbb{R})\simeq(\mathbb{R}^{\times})^{2}\times_{\operatorname{N}}(\mathbb{R}^{\times})^{2}$ are pairs of diagonal matrices with same determinant. The preimage $\widetilde{K}_{\operatorname{T},\infty}=(\phi_{1}\times\phi_{2})^{-1}(\widetilde{K}_{\infty})$ of $\widetilde{K}_{\infty}$ by $\phi_{1}\times\phi_{2}$ (tensored with $\mathbb{R}$) is $\mathbb{R}_{>0}(\pm 1,\pm 1)$. Hence
@@
As in [Bra23], we use that $\langle C_{T},C_{n}(\varphi)\rangle=\langle\gamma_{\mathcal{O}_{1}},T_{n}\gamma_{\mathcal{O}_{2}}\rangle$ to deduce that
@@
is a diagonal restriction of a Hilbert modular form for $\operatorname{SL}_{2}(\mathbb{Q}(\sqrt{D_{1}D_{2}}))$.
@@
    A similar generating series of intersection numbers on a compact shimura curve have been considered by Rickards [Ric22]. It should also follow from the seesaw argument that it is the diagonal restriction of a Hilbert modular form.

### v4

@@ 3.5. Computations of the Hilbert modular form
[…]
Suppose that $E$ is a field. Since $\operatorname{T}$ is of maximal real rank, $S_{3}$ is empty by Proposition 2.6. Recall that $q=|S_{1}|$ by Proposition 2.6. At a place $\sigma\in S_{2}$, the algebra is $E_{\sigma}\simeq\mathbb{C}$ and the space $\mathcal{D}_{\sigma}^{+}$ is a point. So the Kudla-Millson form is of degree 0 and is simply the Gaussian
@@
At a place $\sigma\in S_{1}$, the algebra is $E_{\sigma}\simeq\mathbb{R}\times\mathbb{R}$ and $\mathcal{D}_{\sigma}^{+}\simeq\mathbb{R}_{>0}$. An element $v$ in $E_{\sigma}$ is sent to $(v_{\sigma},v^{\prime}_{\sigma})$ in $E_{\sigma}$. The Kudla-Millson form (see [5]) is given by
@@
for $g_{\infty}\in\operatorname{SL}_{2}(F_{\mathbb{R}})$. Note that since $E$ is a field, any nonzero $v\in E$ must satisfy $v_{\sigma}v_{\sigma}^{\prime}\neq 0$ at any place $\sigma\in S_{1}$. Let $\tau\coloneqq(\tau_{\sigma})_{\sigma}\in\mathscr{H}^{d}$ be the image of $(i,\dots,i)$ by $g_{\infty}$ and write
@@
for $m\in F$. The following Lemma is a classical computation. A similar computation appears in Kudla’s work [18], and even Hecke’s original paper [11].
@@
$\int_{E_{\infty}^{1}}\omega(g_{\infty},t_{\infty})\varphi_{\infty}(v)\chi_{\infty}(t)dt_{\infty}^{\times}=2^{q}\left(y_{1}\cdots y_{d}\right)^{\frac{1}{2}}\left(\prod_{\sigma\in S_{1}}\operatorname{sgn}(v_{\sigma}+v^{\prime}_{\sigma})\right)e^{2i\pi\operatorname{Tr}_{F/\mathbb{Q}}(\tau\alpha m)}$
if $m=\operatorname{N}_{E/F}(v)$ is totally positive, and the integral is 0 otherwise.
@@
$\int_{E_{\sigma}^{1}}\omega(1,{t_{\sigma}})\varphi_{\sigma}(v_{\sigma},v_{\sigma}^{\prime})\chi_{\sigma}(t_{\sigma})dt_{\sigma}^{\times}$ $=\sqrt{|\alpha_{\sigma}|}\int_{\mathbb{R}^{\times}}e^{-\pi|\alpha_{\sigma}|\left((v_{\sigma}{t_{\sigma}}^{-1})^{2}+({t_{\sigma}}v_{\sigma}^{\prime})^{2}\right)}\left(\frac{v_{\sigma}}{{t_{\sigma}}}+{t_{\sigma}}v_{\sigma}^{\prime}\right)\operatorname{sgn}_{\sigma}(t_{\sigma})\frac{d{t_{\sigma}}}{{t_{\sigma}}}$
@@
after the substitution $t_{\sigma}=\sqrt{\left|\frac{v_{\sigma}}{v_{\sigma}^{\prime}}\right|}u$. Note that the assumption that $\chi_{\sigma}(t_{\sigma})=\operatorname{sgn}(t_{\sigma})$ is crucial here, since the integral would be 0 otherwise. After substituting $\beta=u^{2}$ we find that this integral is
@@
$\int_{E_{\infty}^{1}}\omega(1,t_{\infty})\varphi_{\infty}(v)\chi_{\infty}(t)dt_{\infty}^{\times}=2^{q}\left(\prod_{\sigma\in S_{1}}\operatorname{sgn}(v_{\sigma}+v^{\prime}_{\sigma})\right)\prod_{\sigma\in S_{1}\cup S_{2}}e^{-2\pi|\alpha_{\sigma}v_{\sigma}v^{\prime}_{\sigma}|}.$ (3.79)
@@
$\Theta^{T}_{\varphi}(\tau_{1},\cdots,\tau_{d})=\widetilde{\Theta}^{T}_{\varphi}(\tau_{1},\cdots,\tau_{d},t)dt^{\times}\in\Omega^{q}(Y_{T}).$ (3.80)
By (3.75) we have
$\widetilde{\Theta}^{T}_{\varphi}(\tau_{1},\cdots,\tau_{d},t)=(y_{1}\cdots y_{d})^{-\frac{1}{2}}\sum_{v\in E}(\omega(g_{\infty},t)\varphi)(v)$ (3.81)
@@
$\Theta^{T}_{\varphi}(\tau_{1},\cdots,\tau_{d},\chi)$ $=\frac{2^{q}}{\operatorname{vol}(K_{T})}\int_{E^{1}\backslash E^{1}_{\mathbb{A}}}(y_{1}\cdots y_{d})^{-\frac{1}{2}}\sum_{v\in E}(\omega(g_{\infty},t)\varphi)(v)\chi(t)dt^{\times}$
$=\frac{2^{q}}{\operatorname{vol}(K_{T})}(y_{1}\cdots y_{d})^{-\frac{1}{2}}\sum_{v\in E^{1}\backslash E}\int_{E^{1}_{\mathbb{A}}}(\omega(g_{\infty},t)\varphi)(v)\chi(t)dt^{\times}$
@@
Let $\chi$ be a character on a anisotropic maximal $\mathbb{Q}$-torus $\operatorname{T}$ of maximal real rank. Then $\Theta^{V}_{\varphi}(\tau,C_{\chi})$ is a cusp form.
@@
Suppose that $F$ is a field, and $E$ a quadratic field extension. Then the only $v\in E$ for which $Q_{\alpha}(v,v)=\operatorname{N}_{E/F}(v,v)=0$ is $v=0$. Since $v_{\sigma}=v_{\sigma}^{\prime}=0$, we see from (3.5.2) that the constant term $c^{(0)}_{\varphi}(C_{\chi})$ is zero.
@@ 3.6. Generating series of intersection numbers
Let $(V,Q)$ be a quadratic space of even dimension $2d$ over $\mathbb{Q}$. Let $L$ be an even integral lattice and $\varphi_{f}=\mathbf{1}_{\widehat{L}}\in\mathcal{S}(V_{\mathbb{A}_{f}})$ its characteristic function. We describe the geometric features of the Kudla-Millson form, that allows us to have a nice geometric interpretation of Kudla-Millson lift $\Theta_{\varphi}(\tau,C)$ defined in (3.3.1).
@@
[…]
@@
[…]
@@
$\Theta_{\varphi}(\tau)=y^{-\frac{p+q}{2}}\sum_{v\in V}(\omega(g_{\tau})\varphi)(v)=\sum_{n\in\mathbb{Q}}\Theta_{\varphi}^{(n)}(\tau)q^{n}$ (3.94)
@@
$\Theta_{\varphi}^{(n)}(\tau)\coloneqq\sum_{\begin{subarray}{c}v\in V\\ Q(v,v)=2n\end{subarray}}\varphi^{0}(\sqrt{y}v)\varphi_{f}(v)=\sum_{\begin{subarray}{c}v\in L\\ Q(v,v)=2n\end{subarray}}\varphi^{0}(\sqrt{y}v).$ (3.95)
The form $\Theta_{\varphi}^{(n)}(\tau)$ is closed, and for positive $n$ it represents a Poincaré dual of the special cycle $C_{n}(\varphi)$ in $H^{q}(Y,\mathbb{C})$. For negative $n$, the form is exact. When $V$ is anisotropic, the constant term is $\varphi_{0}(0)\varphi_{f}(0)$ where $\varphi_{0}(0)$ represents the Euler class of the tautological bundle over $\mathcal{D}^{+}$.
@@
$\Theta^{V}_{\varphi}(\tau,C)=\int_{C}\Theta^{V}_{\varphi}(\tau)$ (3.96)
@@
$\int_{C}\Theta^{V}_{\varphi}(\tau)=c^{0}_{\varphi}(C)+\sum_{n=1}^{\infty}\int_{C}\Theta_{\varphi}^{(n)}(\tau)q^{n}=c^{0}_{\varphi}(C)+\sum_{n\in\mathbb{Q}_{>0}}\langle C,C_{n}(\varphi)\rangle q^{n}.$ (3.97)
When the cycles $C$ and $C_{n}(\varphi)$ intersect transversely, then $\langle C,C_{n}(\varphi)\rangle$ is the signed intersection number.
@@
[…]
@@ 4. Example of biquadratic fields
We consider the setting of a split quadratic space of signature $(2,2)$. The setting is similar to [4], but we consider the integral over a product of compact geodesics instead of the the product of a compact geodesic with $\gamma_{\infty}$.
@@
The connected component $\operatorname{GSpin}_{V}(\mathbb{R})^{+}=\operatorname{GL}_{2}(\mathbb{R})^{+}\times_{\det}\operatorname{GL}_{2}(\mathbb{R})^{+}$ consists of pairs of matrices with (same) positive determinant. It acts transitively on $\mathscr{H}\times\mathscr{H}$ by Möbius transformations in both factors. Let $\widetilde{K}_{\infty}\subset\operatorname{GSpin}_{V}(\mathbb{R})^{+}$ be the stabilizer of a point in $\mathscr{H}\times\mathscr{H}$, so that $\mathscr{H}\times\mathscr{H}\simeq\operatorname{GSpin}_{V}(\mathbb{R})^{+}/\widetilde{K}_{\infty}$. The stabilizer of $(i,i)$ is $\mathbb{R}_{>0}(\operatorname{SO}(2)\times\operatorname{SO}(2))$. We can extend the Weil representation from the pair $\operatorname{SO}_{V}(\mathbb{Q}_{v})\times\operatorname{SL}_{2}(\mathbb{Q}_{v})$ to the pair $\operatorname{GSpin}_{V}(\mathbb{Q}_{v})\times\operatorname{SL}_{2}(\mathbb{Q}_{v})$ by
@@ 4.1. Locally symmetric space
For the locally symmetric spaces and the special cycles in this setting, we refer to [4] for more details. Let $M_{0}(p)\subset\operatorname{Mat}_{2}(\mathbb{Q})$ be the lattice
@@
[…]
$K_{0}(p)=\left\{\left.\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{GL}_{2}(\widehat{\mathbb{Z}})\ \right|\ c\in p\widehat{\mathbb{Z}}\right\}$. (4.10)
@@
by the strong approximation $\operatorname{GL}_{2}(\mathbb{A}_{f})=\operatorname{GL}_{2}(\mathbb{Q})^{+}K_{0}(p)$.
@@
[…]
@@ 4.2. Maximal torus in $\operatorname{GSpin}_{V}$
Let $L=\mathbb{Q}(\sqrt{D})$ be a real quadratic field of fundamental discriminant $D$ and suppose that $p$ is split in $L$. Let $\mathcal{O}_{L}$ be the ring of integers. Let us fix an integer $r\in\mathbb{Z}$ such that $r^{2}\equiv D\pmod{4p}$. A form $a,b,c]=ax^{2}+bxy+cy^{2}$ of squarefree discriminant $D$ is called a Heegner form at $p$ if it satisfies $a\equiv 0\pmod{p}$ and $b\equiv r\pmod{p}$. Let $(u,v)$ be a positive fundamental solution to the Pell equation $u^{2}-Dv^{2}=1$. To every primitive Heegner form $[a,b,c]$ at $p$ of discriminant $D=b^{2}-4ac$ we can associate the matrix
@@
[…]
$1\xrightarrow{\ \ \ \ \ }\mathbb{Q}^{\times}\xrightarrow{\ \ \ \ \ }\widetilde{\operatorname{T}}(\mathbb{Q})\xrightarrow{\ \ \rho\ \ }\operatorname{T}(\mathbb{Q})\xrightarrow{\ \ \ \ \ }1.$ (4.17)
Let $L_{1}$ and $L_{2}$ be two real quadratic fields with distinct discriminants $D_{1}$ and $D_{2}$. Suppose that $p$ is split in both fields. Let $N_{1},r_{1},1]$ and $N_{2},r_{2},1]$ be the two principal [^3] It represents the unit in the narrow class group. Heegner forms where $N_{i}\coloneqq\frac{r_{i}^{2}-D_{i}}{4}$. Let $\phi_{i}\colon L_{i}\lhook\joinrel\longrightarrow\operatorname{Mat}_{2}(\mathbb{Q})$ be the two associated optimal embeddings, given by
@@
We can then map $\operatorname{GSpin}_{V}(\mathbb{Q})$ in $\operatorname{End}_{\mathbb{Q}}(\operatorname{Mat}_{2}(\mathbb{Q}))$ by sending $(g_{1},g_{2})$ to $\det(g_{2})g_{1}\otimes g_{2}$. The map is compatible with the actions of $\operatorname{GSpin}_{V}(\mathbb{Q})$ and $\operatorname{End}_{\mathbb{Q}}(V)$ on $V$.
@@
[…]
@@
The étale algebra of endomorphisms commuting with $J$ is $E=L_{1}\otimes L_{2}$. Note that when restricted to $L_{i}$ the involution $x^{*}$ acts like the Galois involution $x\mapsto x^{\prime}$, i.e. we have $\phi(x^{\prime})=\phi(x)^{*}$. Hence, the involution $\epsilon_{Q}$ restricted to $E$ is $\epsilon_{Q}(x_{1}\otimes x_{2})=x_{1}^{\prime}\otimes x_{2}^{\prime}$ where $x_{i}^{\prime}$ is the Galois conjugate of $x_{i}$. It follows that the étale algebra is $E\simeq\mathbb{Q}(\sqrt{D_{1}},\sqrt{D_{2}})$ with the involution sending $\sqrt{D_{i}}$ to $\sqrt{D_{i}}$. ∎
@@
$\mathbb{Q}(\sqrt{D_{1}},\sqrt{D_{2}})^{1}\simeq\operatorname{T}(\mathbb{Q})\simeq\rho\left(L_{1}^{\times}\times_{\operatorname{N}}L_{2}^{\times}\right).$ (4.25)
@@
[…]
@@
As in [4], we use that $\langle C_{T},C_{n}(\varphi)\rangle=\langle\gamma_{\mathcal{O}_{1}},T_{n}\gamma_{\mathcal{O}_{2}}\rangle$ to deduce that
@@
is a diagonal restriction of a Hilbert modular form for a subgroup of $\operatorname{SL}_{2}(\mathbb{Q}(\sqrt{D_{1}D_{2}}))$.
@@
    A similar generating series of intersection numbers on a compact Shimura curve have been considered by Rickards [22]. It should also follow from the seesaw argument that it is the diagonal restriction of a Hilbert modular form.